# MtgrBoost: Boosting Large-scale Generative Recommendation Models in Meituan

### v1

@@ 5.1 Sequence Balancing
       $k\leftarrow$ binary search in $S$ for closest value to $N$
@@
Figure 10: Running example of dynamic sequence batching. The number indicates the number of tokens in a sample.
Dynamic Sequence Batching Algorithm 1 illustrates MTGRBoost’s dynamic sequence batching mechanism, where each GPU maintains a buffer $\mathcal{Q}$ for storing sequence samples retrieved from hive table chunks $C_{i}\in\mathcal{C}$. We define target token count $N$ as the product of average sequence length (i.e., 600) and batch size. The system first computes token counts per sample in the sequence buffer and calculates cumulative sums $S$. A binary search algorithm then identifies the optimal partition point $k$ where the cumulative sum most closely approaches the target token count. We finally obtain a balanced batch $\mathcal{B}$ by outputting the first $k$ sequences in the buffer. When total buffer tokens fall below the target, remaining samples merge into subsequent buffer. This process iterates until all hive table chunk samples are consumed.
In data parallelism, gradients are typically averaged using All-Reduce operations. However, with dynamic sequence batching, directly averaging gradients may introduce biases, as each device computes gradients based on a varying number of samples. To address this issue, we implement All-to-all communication to synchronize batch sizes across devices, followed by performing a weighted average of gradients proportional to their respective batch sizes. This mechanism ensures numerical correctness in both loss computation and gradient update processes.
Discussions. We observe two critical advantages of dynamic sequence batching in distributed training systems. First, device load balancing through constraining token counts. Second, optimized hardware utilization via dynamic batch sizes. As shown in Figure 10, GPU 0 requires 500 batches to reach the threshold while GPU 1 achieves it with 200 batches. Fixed size strategy requires conservative batch size configurations to prevent out-of-memory risks from clusters of extremely long sequences. In contrast, dynamic batching maximizes device memory utilization by loading batches near device memory limits. While more complex load balancing methods exist, including sequence packing based on length distributions [17, 19] or dynamic programming-based parallelism strategies [20], we ultimately exclude these methods from our MTGRBoost. This selection stems from the empirical observation that our lightweight dynamic sequence balancing achieves near-optimal workload balancing under current training conditions, rendering additional complexity unnecessary.
5.2 Implementations
Checkpoint Resuming. The core objective of checkpoint is to achieve persistent model preservation during training, enabling rapid recovery to previous training states. MTGRBoost not only supports conventional checkpoint resuming but also facilitates preservation and loading in dynamic distributed environments (e.g., saving on 8 GPUs but loading on 16 GPUs). Existing systems address this challenge by saving model parameters from all devices in a single checkpoint, then redistributing parameters based on new device quantities during loading. However, this approach suffers from a critical flaw: each device must scan the entire checkpoint, resulting in redundant read operations. In contrast, MTGRBoost implements a novel approach where each device independently preserves its own checkpoint. During loading, new devices locate required checkpoint files through modulo operations. For instance, when loading checkpoints saved from 8 GPUs onto 16 GPUs, both GPU 0 and GPU 8 load parameters from the checkpoint saved on the original GPU 0. This design is grounded in the insight that distributed cluster scaling typically follows powers of two.
@@ 5.2 Implementations
Gradient Accumulation. MTGRBoost enhances training stability by accumulating gradients across multiple training batches to mitigate large gradient variances caused by small batch sizes. For sparse embedding parameters, we first record activated embedding IDs and their corresponding gradient values within each batch. These gradients from identical IDs across multiple batches are accumulated and then updated collectively. Furthermore, we avoid full parameter updates for sparse embeddings, instead selectively updating only activated parts. This design simultaneously reduces memory waste and improves update efficiency. For smaller dense models, we also implement gradient accumulation followed by full parameter updates.
@@
Table 1: Model hyperparameters in GRM. Complexity denotes computational complexity, and the unit is FLOPs.
@@
6 Experimental Evaluation
6.1 Experiment Settings
Model Setup. For GRM’s dense model, we scale model architecture based on computational complexity, resulting in GRM 4G and GRM 110G variants, where “G” corresponds to Giga Floating-Point Operations (GFLOPs) required per forward pass. For sparse models, we use the embedding dimension factor to expand the embedding tables, including 1D, 8D, and 64D configurations. The baseline 1D configuration chooses the widely adopted embedding dimensions [36, 41, 25], typically ranging from 32 to 128 (with variations across different features). Accordingly, 64D represents a 64$\times$ expansion of all embedding table dimensions relative to this 1D baseline. We adjust the embedding dimension factor in experiments including time decomposition, lookup deduplication, dynamic table, and scalability due to these experiments are sensitive to embedding dimension, and keep a default value of 1 in other experiments. The detailed model hyperparameters are provided in Table 1.
Baselines. We compare MTGRBoost with TorchRec, the state-of-the-art GRM training framework proposed by Meta [10]. Actually, our initial design is developed based on TorchRec, and we further implement MTGRBoost with the proposed optimizations. We conduct experimental comparisons against TorchRec regarding both model accuracy and efficiency. In the ablation experiments, we investigate the effects of our designs in MTGRBoost. We exclude alternative baselines as TorchRec is the only PyTorch-native training framework specifically optimized for PyTorch-based GRMs. Although other training systems such as TensorFlow are widely used in the industry, their framework currently cannot support GRM training.
Dataset and Implementation. We conduct experiments on 90 days of user logs from Meituan, with daily log data exceeding 400 million sequences. The average sequence length is 600, reaching a maximum length of 3,000. All models are trained from scratch using the same data. To optimize sparse and dense features, we employ the Adam [16] optimizer.
Evaluation Metrics. We use Group AUC to evaluate the accuracy of the model on CTR and CTCVR tasks and system throughput to evaluate the training efficiency. GAUC calculates the AUC metric by grouping at the user level, which can better reflect the actual performance of the recommendation model compared to conventional global AUC.
6.2 Comparisons to Baselines
Figure 11: The GAUC results of CTR and CTCVR on GRM 4G 1D for TorchRec and MTGRBoost, respectively.
Figure 12: Time decomposition for GRM 4G 1D and GRM 100G 64D in TorchRec and MTGRBoost, respectively.
Accuracy. Figure 11 presents the CTR and CTCVR GAUC results obtained from training a generative recommendation model using MTGRBoost and TorchRec. The GRM models in two systems exhibit rapid GAUC growth during the initial 40,000 training steps, followed by gradual improvement over the subsequent 30,000 steps before stabilizing. This progression demonstrates that both MTGRBoost and TorchRec ensures both model correctness and training stability.
Time Decomposition. Figure 12 presents the time decomposition of GRM 4G 1D and GRM 50G 64D across 100 cumulative training steps, detailing execution times for embedding lookup, forward, and backward phases. We can see that MTGRBoost achieves shorter execution times across all phases compared to TorchRec. The efficiency improvements derived from sequence balancing in dense model computations grow progressively more pronounced with increasing model complexity. Notably, expanding embedding table dimensions induces significant temporal increases in both lookup and backward phases, primarily stemming from embedding communication overhead and sparse parameter update latency. However, our implementation of automatic table merging and lookup deduplication substantially enhances execution efficiency in these critical phases.
Figure 13: Ablation study for GRM 4G 1D and GRM 110G 1D in MTGRBoost. Compared with the baseline, MTGRBoost achieves a 1.60×1.60$\times$1.60 × to 2.44×2.$4$ throughput improvement.
Figure 14: Throughput comparison for disabling and enabling sequence balancing on GRM 4G 1D and 110G 1D. We gradually scale the system from 8 GPUs to 64 GPUs. We report the gain values of sequence balancing at the top of figure.
6.3 Evaluation of Individual Designs
Ablation Study. Figure 13 presents an ablation study under 4G and 10G complexity settings by incrementally integrating different techniques in MTGRBoost. Starting with merging tables, we consolidate ID communication, embedding communication, and lookup operators from multiple feature tables into a single operation. Subsequently, a two-stage deduplication process is applied to feature IDs prior to embedding lookups. Finally, we introduce the sequence balancing strategy to optimize device load distribution, activating all MTGRBoost’s components.
The experimental results demonstrate consistent improvements across all proposed designs, indicating that every component in MTGRBoost is critical for enhancing overall system efficiency. Furthermore, we observe that performance gains intensify with higher computational complexity. This trend reveals a positive correlation between MTGRBoost ’s efficiency improvements and computational complexity.
Table 2: Batch sizes and average GPU memory utilization for three models with sequence balancing disabled versus enabled. Values before the arrow denotes fixed batch sizes with disabled state, while those after the arrow represent average ones over 1,000 steps with enabled state. Numbers in parentheses denote improvements in memory utilization.
Model Batch size GPU memory utilization
GRM 4G 1D 480 → 496 86.3% → 95.7% (+9.4)
GRM 110G 1D 80 → 116 75.3% → 90.3% (+15.0)
Figure 15: Original sequence vs. dynamic sequence batching. Each box denotes the minimum and maximum total token counts for GRM 4G 1D across 8 GPUs in one training step.
Sequence Balancing Evaluation. We further evaluate the proposed sequence balancing. We conduct experiments with GRM 4G and 110G. Starting with 8 GPUs, then we double the GPU count up to 64. To ensure fairness, batch sizes are adjusted to maintain near-full GPU utilization under both sequence balancing enabled and disabled conditions. Table 2 reports batch sizes and average GPU memory utilization, while Figure 14 presents throughput results with and without sequence balancing, yielding five key observations.
  • $\bullet$
    Enabling sequence balancing in MTGRBoost achieves significant throughput improvements across all model variants and GPU configurations. Specifically, GRM 4G, and 110G exhibit average throughput improvements of 4.4%, and 26.5%, respectively, with GRM 110G on 64 GPUs showing a peak improvement of 33.5%.
  • $\bullet$
    For a given model, throughput improvements scale with GPU count under sequence balancing. This is because that each step is constrained by the slowest device in synchronous training architectures. As the number of GPUs increases, the likelihood of encountering long sequences that induce computational delays rises.
  • $\bullet$
    Throughput gains correlate positively with computational complexity. Since sequence processing FLOPs scale quadratically with token hidden dimensions, higher computational complexity amplifies load imbalance, which makes sequence balancing increasingly critical.
  • $\bullet$
    Figure 15 illustrates reduced token count variance across GPUs after enabling sequence balancing, which stabilizes token counts at approximately 76,000 per device, effectively mitigating load imbalance.
  • $\bullet$
    Table 2 shows that dynamic sequence batching can maximize GPU memory utilization via dynamic batch size regulation. In contrast, raw sequences would necessitate smaller batches to avoid out-of-memory errors from extreme sequence lengths, thereby degrading system efficiency.
Figure 16: Throughput comparisons of two-stage ID deduplication for GRM 4G 1D (upper) and GRM 4G 64D (bottom). The improvements of the current strategy compared to the previous one are reported above the bars.
The Effect of Two-stage ID Deduplication. To further evaluate the impact of two-stage ID deduplication on system efficiency, we fix the computational complexity at 4GFlOPs and conduct experiments with two embedding dimension factors of 1D and 64D. We employ four deduplication strategies: (a) w/o unique (i.e., no deduplication), (b) Comm. unique (i.e., first-stage deduplication only), (c) Lookup unique (i.e., second-stage deduplication only), and (d) Two-stage unique (i.e., both stages). Starting with 16 GPUs, we incrementally increase the GPU count and report throughput results in Figure 16. We get the following observations.
First, the two-stage approach achieves a 1.1$\times$–3.7$\times$ throughput improvement compared to the baseline (a), with gains amplifying at larger GPU scales. This finding highlights the effectiveness of two-stage deduplication in reducing ID communication, embedding communication, and lookup latency. Second, among single-stage strategies, “Comm. unique” outperforms “Lookup unique”. As embedding communication represents the primary latency source, whereas hash table lookups are inherently fast. Finally, deduplication benefits correlate with embedding dimensions: higher-dimensional embeddings amplify network bandwidth demands during cross-node transfers, which makes ID deduplication more important for reducing communication volume.
Table 3: Throughput comparison for MCH and MTGRBoost. OOM denotes out of memory and Gain denotes the throughput improvement. “-” indicates that the comparison cannot be made due to OOM problems.
Complexity Dim. Factor MCH MTGRBoost Gain
4G 1D 392,731 579,649 47.59%
8D 260,590 438,190 68.2%
64D 80,857 155,832 92.7%
110G 1D 44,719 68,993 54.28%
8D 28,329 55,929 97.4%
64D OOM 38,575 -
The Effect of Dynamic Embedding Table. To evaluate the effectiveness of the proposed dynamic embedding table in handling dynamic IDs, we conduct comparative experiments using Managed Collision Handling (MCH) as the baseline. MCH is designed by TorchRec to resolve changeable feature ID by maintaining a fixed-size mapping table to remap original IDs into a continuous space. It employs binary search for efficient ID localization and activates an eviction mechanism to update ID mappings when a threshold is reached.
Experimental results are shown in Table 3. Using a model with fixed computational complexity, we assess the throughput of both two techniques in handling embedding tables of varying scales by adjusting the embedding dimension factor. Our findings reveal that the dynamic embedding table achieves 1.47$\times$–2.22$\times$ throughput improvement than MCH. This improvement stems from the grouped parallel probing technique, which accelerates hash collision resolution via multi-threaded parallel probing. Furthermore, MCH encounters OOM issues at larger embedding dimensions due to pre-allocated memory for all tables. In contrast, the hash table requires substantially less initial memory and allocates additional resources only when thresholds are exceeded.
Figure 17: Comparison of scalability by (a) computational complexity and (b) embedding dimension. We use 1 node, 8 GPUs configuration as the baseline for speedup ratio. The specific speedup ratio values are marked near the broken line.
6.4 Scalability
We assess the system’s scalability by examining the deviation between actual speedup ratios and ideal linear scaling. Two factors are evaluated: computational complexity and embedding dimensions. (1) With the embedding dimension fixed at 1D, we compare speedup ratios under computational complexities of 4G and 110G. (2) With computational complexity fixed at 4G, we compare speedup ratios for embedding dimensions of 2D and 64D. We use a 1 node, 8 GPUs configuration as the baseline and progressively scale to 16 nodes (in total 128 GPUs). Experimental results in Figure 17 reveal three key findings.
First, all systems exhibit sublinear scaling as GPU count increases, which attributes to communication overhead introduced by parallelism. However, MTGRBoost achieves 62.75%–78.5% of the ideal speedup ratio at 128 GPUs, which demonstrates its robust scalability. Second, scaling computational complexity by 27.5$\times$ (GRM 4G 1D vs. 110G 1D) and embedding dimensions by 32$\times$ (GRM 4G 2D vs. 4G 64D) results in minor speedup degradation. This stability arises from sequence balancing mitigating load imbalance under increased complexity, while merged tables and two-stage ID unique operation reduce lookup and communication latency. Third, embedding dimensions impact speedup more significantly than computational complexity. Since the latency of recommendation system is dominated by sparse embedding communication, computation, and updates, making scaling embedding dimensions more likely to affect performance.
7 Related Work
Systems for DRM training. Deep Recommendation Models (DRMs) combine sparse embeddings and dense neural networks to model each user-item individually. As DRMs are widely used, many systems are developed to accelerate its training [21, 13, 40, 23, 14, 6, 5, 46, 7, 32, 18, 38]. Specifically, Persia [21] decouples sparse embedding computation from dense model processing, training the embeddings asynchronously to improve throughput and the dense models synchronously to ensure accuracy. PICASSO [45] leverages the patterns of model architecture and data distribution to accelerate execution with packing, interleaving, and caching optimizations. XDL [13] compresses the pair-wise communication among GPUs into a tree structure to accelerate embedding lookup and update. ScaleFreeCTR [7] leverages host memory to store massive embedding tables and employs GPU-optimized synchronization to update the embeddings. Zion [32] and RecSpeed [18] mitigate the I/O bottlenecks for embedding communication by deploying additional NICs. HugeCTR [38] tailors DRM training for NVIDIA DGX-1/DGX-2 platforms by optimizing embedding communication on NV-Switch interconnects. These systems are usually designed according to specific assumptions on the access patterns of the sparse embeddings (e.g., skewed towards the hot embeddings). We also plan to incorporate their optimizations after carefully analyzing the embedding access patterns in Meituan data. Moreover, their TensorFlow-based implementations require substantial efforts to migrate to PyTorch environments.
Systems for Transformer. To train Transformer efficiently, the focuses are attention computation and parallel strategies [29, 35, 3, 22, 12, 9, 30, 28, 20, 39, 27]. FlashAttention [3] reduces the accesses to slow GPU global memory with tiled computation and online recomputation fusion. FlexFlow [22] and OptCNN [12] automatically search the optimal parallel strategy for model training by distributing model computation and parameters among the GPUs in different ways. GPipe [9] employs pipeline parallelism by splitting each mini-batches into micro-batches to interleave the computation and communication across the micro-batches. Megatron-LM [30] and DeepSpeed [28] propose hybrid parallel strategies for large-scale model training by partitioning the model and data across the GPUs. Hydraulis [20] adopts dynamic parallelism for each iteration according to the sequence length distribution to achieve workload balance. MTGRBoost enjoys existing attention kernel optimizations for efficient computation but simple data parallel suffices for GRMs since their dense models are usually small. Moreover, MTGRBoost optimizes the processing of sparse embeddings, which are not present in most Transformer-based models.
8 Conclusion
In this paper, we propose MTGRBoost, a distributed training system addressing scalability and efficiency challenges in industrial generative recommendation models (GRMs). For sparse model, we introduce dynamic embedding tables for changeable feature IDs, automated table merging and two-stage ID deduplication for embedding optimizations. For dense model, dynamic sequence balancing alleviates the workload imbalances. A suite of other techniques are integrated to ensure system availability. Experimental results demonstrate substantial improvements in training efficiency while maintaining model accuracy. MTGRBoost provides a flexible and efficient solution for industrial applications, paving the way for broader adoption of GRM in recommendation systems.
Appendix A Proof
Lemma 1
Proof 1
$S$ contains no factor of 2 (since $S$ is odd), while $M$ has prime factorization $2^{n}$. Thus, $S$ and $M$ share no common prime factors, proving $\mathrm{gcd}(S,M)=1$.
Proof 2
Assume there exist $t_{1},t_{2}\!\in\![0,M\!-\!1]$, $t_{1}\!\neq\!t_{2}$ such that:
$h_{t_{1}}\equiv h_{t_{2}},$ (9)
This implies:
$(t_{1}-t_{2})\cdot S\equiv 0.$ (10)
$t_{1}-t_{2}\equiv 0.$ (11)
Proof 3
$\{h_{t}\}_{t=0}^{M-1}=\{0,1,2,...,M-1\}.$ (12)

### v2

@@ B.4. Running Example for Dynamic Offset
 $k\leftarrow$ binary search in $S$ for closest value to NN
@@
B.5. Algorithm for Sequence Balancing
Algorithm 1 depicts MTGenRec’s dynamic sequence batching. Each GPU maintains a sequence buffer $\mathcal{Q}$ populated from hive table chunks $C_{i}\in\mathcal{C}$. The target token count $N$ equals the batch size multiplied by the average sequence length. Token counts are aggregated into a cumulative sum $S$ over samples in $\mathcal{Q}$. A binary search finds the partition index $k$ where $S(k)$ best approximates $N$. The first $k$ sequences in $\mathcal{Q}$ form the output batch $\mathcal{B}$. Under-filled buffers accumulate samples until processing exhausts all hive chunk data.
Figure 16. An example of offsets in static table and hash table.
Appendix C Implementation Details for MTGenRec
Checkpoint Resuming. The core objective of checkpoint is to achieve persistent model preservation during training, enabling rapid recovery to previous training states. MTGenRec not only supports conventional checkpoint resuming but also facilitates preservation and loading in dynamic distributed environments (e.g., saving on 8 GPUs but loading on 16 GPUs). Existing systems address this challenge by saving model parameters from all devices in a single checkpoint, then redistributing parameters based on new device quantities during loading. However, this approach suffers from a critical flaw: each device must scan the entire checkpoint, resulting in redundant read operations. In contrast, MTGenRec implements a novel approach where each device independently preserves its own checkpoint. During loading, new devices locate required checkpoint files through modulo operations. For instance, when loading checkpoints saved from 8 GPUs onto 16 GPUs, both GPU 0 and GPU 8 load parameters from the checkpoint saved on the original GPU 0. This design is grounded in the insight that distributed cluster scaling typically follows powers of two.
@@ Appendix C Implementation Details for MTGenRec
Gradient Accumulation. MTGenRec enhances training stability by accumulating gradients across multiple training batches to mitigate large gradient variances caused by small batch sizes. For sparse embedding parameters, we first record activated embedding IDs and their corresponding gradient values within each batch. These gradients from identical IDs across multiple batches are accumulated and then updated collectively. Furthermore, we avoid full parameter updates for sparse embeddings, instead selectively updating only activated parts. This design simultaneously reduces memory waste and improves update efficiency. For smaller dense models, we also implement gradient accumulation followed by full parameter updates.
@@
Appendix D Experiment Details
D.1. Datasets
Public datasets widely use independent ID and attribute features, where cross features are seldom introduced. However, cross features demonstrate importance in real-world application. Cross features are an important category of features in our scenario. They are often meticulously hand-crafted, including interactions like user-item, user and higher-level categories, item and spatio-temporal information. To compensate the absence of cross feature in public datasets, we construct a training dataset based on logs from the real industrial-scale recommendation system in Meituan. The detailed dataset statistics are reported in Table 5. Unlike public datasets, our real dataset contains a richer cross features set and longer user behavior sequences. In addition, the volume of our dataset is large, allowing complex models to achieve more adequate convergence during training. For the offline experiments, we collect data over a $10$-day period. For the online experiments, in order to compare with the DLRM baseline that has been trained for over $2$ years, we constructed a longer-term dataset for the experiments, using data spanning more than $6$ months.
D.2. Model Hyperparameters
We use two models: small and large GRM for all the experimental evaluations. The detailed model hyperparameters are provided in Table 6.
Table 5. Dataset Statistics
Dataset #Users #Items #Exposure #Click #Purchases
Train 0.21 billion 4,302,391 23.74 billion 1.08 billion 0.18 billion
Test 3,021,198 3,141,997 76,855,608 4,545,386 769,534
Table 6. Model hyperparameters in GRM. Complexity denotes computational complexity, and the unit is FLOPs.